# Shift-cyclicity in Analytic Function Spaces

### v1

@@ Abstract.
In this survey, we consider Banach spaces of analytic functions in one and several complex variables for which: (i) polynomials are dense, (ii) point-evaluations on the domain are bounded linear functionals, and (iii) the shift operators are bounded for each variable. We discuss the problem of determining the shift-cyclic functions in such a space, i.e., functions whose polynomial multiples form a dense subspace. The problem of determining shift-cyclic functions in certain analytic function spaces is known to be intimately connected to some deep problems in other areas of mathematics, such as the dilation completeness problem and even the Riemann hypothesis.
@@ 1. Introduction
If $\mathcal{X}$ is finite dimensional, then the eigenspaces corresponding to $T$ are clearly all $T$-invariant. Now, if $\mathcal{X}$ is a complex vector space, then we know that such eigenspaces always exist and therefore, any linear map on a finite dimensional complex Banach space has a non-trivial invariant subspace. Thus, it is a natural question to ask if one can similarly always find closed non-trivial invariant subspaces for any given $T\in\mathcal{L}(\mathcal{X})$ even when $\mathcal{X}$ is infinite dimensional. This is the famous invariant subspace problem (ISP) [^1] ISP was inspired by unpublished works of Beurling and von Neumann circa mid-$20^{\text{th}}$ century. Halmos [Hal63] then popularized the ISP through expository writing., which has been resolved in specific cases [[, see]]AH11, AS54, Atz83, Bea85, BR66, Enf75, Hal66, Lom73, Rea85, Sli08, but remains open – as of the writing of this article – for the case when $\mathcal{X}$ is a separable Hilbert space.
@@
Evidently, determining the cyclicity of a vector is challenging in general. For finite dimensional spaces, however, determining whether an operator is cyclic is a fairly standard exercise in linear algebra based on the following result [[, see]the second corollary to Theorem 3 in Section 7.2]HK73.
@@
If $T$ is cyclic on a finite dimensional space $\mathcal{X}$, then there exist computational algorithms to generate the $T$-cyclic vectors as well [[, see]Sections 5 and 6]AC97. This is not true in general if $\mathcal{X}$ is infinite dimensional, and we typically need to employ other tools – specific to the space – in order to determine cyclic vectors. We showcase this with a concrete example in the next subsection.
@@ 1.1. Smirnov–Beurling theorem
Let $\mathbb{D}$ denote the unit disk in the complex plane $\mathbb{C}$, and let $\operatorname{Hol}(\mathbb{D})$ be the vector space of holomorphic functions on $\mathbb{D}$. Now, we define
@@
$\left\langle f=\sum_{k\in\mathbb{Z}_{+}}a_{k}z^{k},g=\sum_{k\in\mathbb{Z}_{+}}b_{k}z_{k}\right\rangle_{H^{2}(\mathbb{D})}:=\sum_{k\in\mathbb{Z}_{+}}a_{k}\overline{b_{k}},$
and is unitarily equivalent to $\ell^{2}(\mathbb{Z}_{+})$ via the correspondence $z^{k}\leftrightarrow e^{(k)}$. Under this correspondence, $S_{\ell^{2}}$ becomes the Hardy shift $S_{H^{2}}$ given by
@@
and that $S_{\ell^{2}}$-cyclic vectors correspond to $S_{H^{2}}$-cyclic functions. Smirnov [Smi32] and Beurling [Beu49] independently identified $S_{H^{2}}$-cyclic functions.
@@
(1.3) $-\infty<\log|f(0)|=\int_{0}^{2\pi}\log|f^{*}(e^{i\theta})|\frac{d\theta}{2\pi}.$
@@
$f^{*}(e^{i\theta}):=\lim_{r\to 1^{-}}f(re^{i\theta})$
which is known to exist for a.e. $\theta\in[0,2\pi)$ and for each $f\in H^{2}(\mathbb{D})$ (see Theorem 2.3). A function satisfying (1.3) is called an outer function. Outer functions play a crucial role in understanding cyclicity beyond $H^{2}(\mathbb{D})$. Note that outerness cannot be realized as a property of the power-series coefficients and thus, the Smirnov–Beurling theorem builds an important bridge between operator theory and function theory. It is also important to note that such a result is rare even for general function spaces over $\mathbb{D}$. Nevertheless, the Smirnov–Beurling theorem spearheaded a deeper analysis of the shift operator $S:f\mapsto zf$, its invariant subspaces, and its cyclic vectors.
@@ 1.2. Cyclicity in general
$\tau:=\left\{T_{g}\right\}_{g\in\mathcal{G}}\subset\mathcal{L}(\mathcal{X})$
@@
(1.4) $\chi_{[0,1]}\in\vartheta[\varrho],\text{ where }\varrho(x):=\left\{\frac{1}{x}\right\}$
@@
This theorem is due to Báez-Duarte [Bae03], who strengthened a similar criterion given by Beurling [Beu55], who in turn strengthened a result of his student Nyman from his Ph.D. thesis [Nym50]. Thus, (1.4) is called the Nyman–Beurling–Báez-Duarte criterion (NBBD) for the Riemann hypothesis.
Somewhat surprisingly, all notions of cyclicity mentioned in Example 1.5 are related to each other – albeit loosely. We give a snippet of this relationship in Section 2.3.1. For now, we refer the reader to the following excellent references on the Euler $\zeta$ function, the Riemann hypothesis and NBBD [Bag06, Edw01, HLS97, Nik19, Noo19].
@@ 1.3.1. Analytic function spaces
and the series converges absolutely and uniformly on all compact subsets of $\Omega$ (see [Rud69, Section 1.1.3] and [Rud80, Section 1.2.6]). Thus, we can endow $\operatorname{Hol}(\Omega)$ with the topology of uniform convergence on compact subsets of $\Omega$. Henceforth, $\mathcal{X}\subset\operatorname{Hol}(\Omega)$ will always represent a subspace under this topology, however, we can guarantee that this holds by introducing other properties that $\mathcal{X}$ satisfies. These properties are important to even make sense of the notion of shift-cyclicity.
@@
A vector space $\mathcal{X}\subset\operatorname{Hol}(\Omega)$ is said to be a Banach/Hilbert space of analytic functions if it is a Banach/Hilbert space that satisfies:
@@
Let us say a few words about P1-P3.
@@
    Depending on what the norm is, P2 – along with Fatou’s lemma – typically guarantees that $\mathcal{X}$ is a subspace of $\operatorname{Hol}(\Omega)$ under the topology of uniform convergence on compact subsets of $\Omega$.
@@
    Note that using the closed graph theorem, if P2 and P3 hold, then $S_{j}\in\mathcal{L}(\mathcal{X})$ for each $1\leq j\leq d$ automatically.
@@ 1.3.2. Shift-cyclic functions
is $\mathbb{Z}_{+}^{d}$-compatible. Also, P1 guarantees that the constant function $1$ is $\sigma$-cyclic. We refer to $\sigma$-cyclicity as shift-cyclicity and refer to $\sigma$-cyclic functions as shift-cyclic functions. Also, we write $S[f]$ instead of $\sigma[f]$.
We end this section with a few properties of shift-cyclic functions that will be important for later discussion. Our first result is immediate from definitions.
@@
(1.5) also implies that a shift-cyclic function $f$ is ‘almost invertible’. It is important to note, however, that having $1/f\in\mathcal{X}$ is not sufficient for the shift-cyclicity of $f$ [[, see]Theorem 1.4]BH97. We shall study more about this algebraic nature of shift-cyclicity in Section 4.2.2 below. Let us also mention another useful property that helps us determine the shift-cyclicity of polynomials.
@@
$\lim_{n\to\infty}\|P-P_{n}Pf\|_{\mathcal{X}}=\lim_{n\to\infty}\|P(1-P_{n}f)\|=0.$
@@
The remaining assertions are immediate from the first part. ∎
@@ 2.1. On the unit disk $\mathbb{D}$
We begin by generalizing the space $H^{2}(\mathbb{D})$ introduced in Section 1.1. The motivation to study these spaces comes from a natural property of functions in $\operatorname{Hol}(\mathbb{D})$ that was discovered by Hardy [Har15]. For any $f\in\operatorname{Hol}(\mathbb{D})$, $r<1$ and $0<p<\infty$, we define the radial $p$-means
@@
By the maximum modulus principle, it is clear that $M_{\infty}(r,f)$ is non-decreasing in $r$ for a fixed $f$, and by the Hadamard three-circles theorem [[, see]Theorem 12.8]Rud74, it follows that $\log M_{\infty}(r,f)$ is a convex function of $\log r$. Hardy observed that the same is true for $M_{p}(r,f)$ for all values of $p$ [^2] Hardy mentioned in his paper that the case $p=1$ was suggested to him by H. Bohr (brother of the physicist N. Bohr) and E. Landau more than a year before he published his result..
Theorem 2.1 (Theorem I, [Har15]).
@@
In this section, we describe a few of the many different properties of Hardy space functions. We have cherry-picked only those that will be important to our discussion on shift-cyclicity. The reader is referred to [Dur70, Gar06, Hof07, Nik19, Rud74] to learn more about the Hardy spaces. We start with the most fundamental of these properties. Let us write $\mathbb{T}:=\partial\mathbb{D}$ for the unit circle.
@@
References. This is essentially a result of Fatou [Fat06], who proved a general result about radial limits of Poisson integrals of complex measures supported on $\mathbb{T}$. Fatou’s theorem gives the case $1\leq p\leq\infty$ while the rest follows from a result of F. and R. Nevanlinna [NN22], which states that for any $f\in H^{p}(\mathbb{D})$ with $0<p<\infty$, there exist $\phi_{1},\phi_{2}\in H^{\infty}(\mathbb{D})$ such that $f=\phi_{1}/\phi_{2}$.
For any $f\in H^{p}(\mathbb{D})$, we define (a.e. on $\mathbb{T}$) its boundary function
@@
Boundary functions play an important role in understanding the structure of functions in the Hardy spaces.
2.1.1. Boundary properties
Let us write $\mathcal{Z}(f)$ for the set of zeros in $\mathbb{D}$ of each $f\in\operatorname{Hol}(\mathbb{D})$ (counting multiplicities) and let $m(0)$ be the multiplicity of the zero at 00 (if $0\in\mathcal{Z}(f)$). Riesz [Rie23] gave the following factorization result.
@@ 2.1.1. Boundary properties
$B(z)=z^{m(0)}\prod_{w\in\mathcal{Z}(f)}\frac{|w|}{w}\frac{z-w}{1-\overline{w}z}.$
@@
It can then be checked that $\left\|\cdot\right\|_{p}$ is a norm on $H^{p}(\mathbb{D})$ for $1\leq p\leq\infty$, and that
@@
is a translation-invariant metric on $H^{p}(\mathbb{D})$ for $10<p<$ 1. Among other things, Riesz’s theorems allow us to obtain the following description of Hardy spaces. See [Dur70, Section 3.2] for all the details.
@@ 2.1.2. Smirnov factorization and shift-cyclicity
    $f$ is said to be inner [^3] Beurling coined the terms inner/outer in [Beu49]. Therefore, the terminology ‘Beurling inner/outer’ is quite common in the literature. if $|f|<1$ on $\mathbb{D}$ and $|f^{*}|=1$ a.e. on $\mathbb{T}$.
@@
    $f$ is said to be outer if $\log|f^{*}|\in L^{1}(\mathbb{T})$ and
    (2.2) $-\infty<\log|f(0)|=\int_{0}^{2\pi}\log|f^{*}(e^{i\theta})|\frac{d\theta}{2\pi}.$
@@
Smirnov [Smi32] and Beurling [Beu49] used this factorization property to show that $f\in H^{2}(\mathbb{D})$ is shift-cyclic if and only if it is outer. In fact, Beurling showed that if we consider the inner part of $f$, i.e., $f_{inn}:=Bf_{\sigma}$, then
@@
Moreover, every closed shift-invariant subspace $E\subset H^{2}(\mathbb{D})$ is of the form $\varphi H^{2}(\mathbb{D})$ where $\varphi$ is the greatest common divisor of the inner factors of each $f\in E$ [[, see]Chapter 6 p. 85]Hof07., i.e.,
@@
Beurling’s approach extends to every $1\leq p<\infty$ using $L^{p}$-duality and the fact that $H^{p}(\mathbb{D})$ can be identified as a closed subspace of $L^{p}(\mathbb{T})$ [[, see]Theorem 7.4]Dur70. Lastly, note that we can also make sense of the shift-invariant subspace $S[f]$ for any $f\in H^{p}(\mathbb{D})$ with $10<p<$ 1 by declaring it as the closure of polynomial multiples of $f$. Gamelin [Gam66] then showed that the Smirnov-Beurling theorem extends to the case $10<p<$ 1 as well. We summarize these results in the following theorem.
@@
This theorem was generalized by Lax [Lax59] and Halmos [Hal61] to the case of vector-valued Hardy spaces. Several other generalizations and modifications of this theorem exist that this article is too short to contain.

### v3

@@ Abstract.
In this survey, we consider Banach spaces of analytic functions in one and several complex variables for which: (i) polynomials are dense, (ii) point-evaluations on the domain are bounded linear functionals, and (iii) the shift operator corresponding to each variable is a bounded linear map. We discuss the problem of determining the shift-cyclic functions in such a space, i.e., functions whose polynomial multiples form a dense subspace. This problem is known to be intimately connected to some deep problems in other areas of mathematics, such as the dilation completeness problem and even the Riemann hypothesis.
@@ 1. Introduction
If $\mathcal{X}$ is finite dimensional, then the eigenspaces corresponding to $T$ are clearly all $T$-invariant. Now, if $\mathcal{X}$ is a complex vector space, then we know that such eigenspaces always exist and therefore, any linear map on a finite dimensional complex Banach space has a non-trivial invariant subspace. Thus, it is natural to ask if one can similarly always find closed non-trivial invariant subspaces for any given $T\in\mathcal{L}(\mathcal{X})$ even when $\mathcal{X}$ is infinite dimensional. This is the famous invariant subspace problem (ISP) [^1] ISP was inspired by unpublished works of Beurling and von Neumann circa mid-$20^{\text{th}}$ century. Halmos [45] then popularized the ISP through expository writing., which has been resolved in specific cases (see [5, 7, 8, 12, 17, 36, 46, 68, 88, 99]), but remains open – as of the writing of this article – for the case when $\mathcal{X}$ is a separable Hilbert space.
@@
Evidently, determining the cyclicity of a vector is challenging in general. For finite dimensional spaces, however, determining whether an operator is cyclic is a fairly standard exercise in linear algebra based on the following result (see [53, the second corollary to Theorem 3 in Section 7.2]).
@@
If $T$ is cyclic on a finite dimensional space $\mathcal{X}$, then there exist computational algorithms to generate the $T$-cyclic vectors as well (see [9, Sections 5 and 6]). This is not true in general if $\mathcal{X}$ is infinite dimensional, and we typically need to employ other tools – specific to the space – in order to determine cyclic vectors. We showcase this with a concrete example in the next subsection.
@@ 1.1. Smirnov–Beurling theorem
Let $\mathbb{D}$ denote the unit disk in the complex plane $\mathbb{C}$ and let $\operatorname{Hol}(\mathbb{D})$ be the vector space of holomorphic functions on $\mathbb{D}$. Now, we define
@@
$\left\langle f=\sum_{k\in\mathbb{Z}_{+}}a_{k}z^{k},g=\sum_{k\in\mathbb{Z}_{+}}b_{k}z^{k}\right\rangle_{H^{2}(\mathbb{D})}:=\sum_{k\in\mathbb{Z}_{+}}a_{k}\overline{b_{k}},$
and is unitarily equivalent to $\ell^{2}(\mathbb{Z}_{+})$ via the correspondence $z^{k}\leftrightarrow e^{(k)}$. Under this correspondence, note that $S_{\ell^{2}}$ becomes the Hardy shift
@@
and that $S_{\ell^{2}}$-cyclic vectors correspond to $S_{H^{2}}$-cyclic functions. Smirnov [101] and Beurling [19] independently identified $S_{H^{2}}$-cyclic functions.
@@
(1.3) $f(0)\neq 0\text{ and }\log|f(0)|=\int_{0}^{2\pi}\log|f^{*}(e^{i\theta})|\frac{d\theta}{2\pi}.$
@@
$f^{*}(e^{i\theta}):=\lim_{r\to 1^{-}}f(re^{i\theta}),$
which is known to exist for a.e. $\theta\in[0,2\pi)$ and for each $f\in H^{2}(\mathbb{D})$ (see Theorem 2.3). A function satisfying (1.3) is called an outer function. Outer functions play a crucial role in understanding cyclicity beyond $H^{2}(\mathbb{D})$. Note that outerness cannot be realized as a property of the power-series coefficients and thus, the Smirnov–Beurling theorem exhibits an interesting overlap between operator theory and function theory. It is also important to note that such a result is rare even for general function spaces over $\mathbb{D}$. Nevertheless, the Smirnov–Beurling theorem spearheaded a deeper analysis of the shift operator $S:f\mapsto zf$, its invariant subspaces and its cyclic vectors.
@@ 1.2. Cyclicity in general
$\tau=\left\{T_{g}\right\}_{g\in\mathcal{G}}\subset\mathcal{L}(\mathcal{X})$
@@
(1.4) $\chi_{[0,1]}\in\vartheta[\varrho],$
where $\varrho\in L^{2}(0,\infty)$ is the function
$\varrho(x):=\left\{\frac{1}{x}\right\}.$
@@
This theorem is due to Báez-Duarte [10], who strengthened a similar criterion given by Beurling [20], who in turn strengthened a result of his student Nyman from his Ph.D. thesis [85]. Thus, (1.4) is called the Nyman–Beurling–Báez-Duarte criterion (NBBD) for the Riemann hypothesis.
Somewhat surprisingly, all notions of cyclicity mentioned in Example 1.5 are related to each other – albeit loosely. We give a snippet of this relationship in Section 2.3.1. For now, we refer the reader to the following excellent references on the Euler $\zeta$ function, the Riemann hypothesis and NBBD [11, 34, 50, 83, 84].
@@ 1.3.1. Analytic function spaces
and the series converges absolutely and uniformly on all compact subsets of $\Omega$ (see [93, Section 1.1.3] and [95, Section 1.2.6]). Thus, we can endow $\operatorname{Hol}(\Omega)$ with the topology of uniform convergence on compact subsets of $\Omega$. Henceforth, $\mathcal{X}\subset\operatorname{Hol}(\Omega)$ will always represent a linear subspace with a topology such that convergence in $\mathcal{X}$ implies convergence in the topology of uniform convergence on compact subsets. In fact, we can guarantee that this holds by introducing other properties that $\mathcal{X}$ satisfies. These properties are important to even make sense of the notion of shift-cyclicity.
@@
A vector space $\mathcal{X}\subset\operatorname{Hol}(\Omega)$ is said to be a Banach/Hilbert space of analytic functions if it is a Banach/Hilbert space with norm $\|\cdot\|_{\mathcal{X}}$ that satisfies:
@@
Let us mention a few words about the properties P1-P3.
@@
    Depending on what the norm is, P2 – along with Fatou’s lemma – typically guarantees that convergence in $\mathcal{X}$ implies uniform convergence on compact subsets of $\Omega$.
@@
    Note that, using the closed graph theorem, if P2 and P3 hold then $S_{j}\in\mathcal{L}(\mathcal{X})$ for each $1\leq j\leq d$.
@@ 1.3.2. Shift-cyclic functions
is $\mathbb{Z}_{+}^{d}$-compatible. Also, P1 guarantees that the constant function $1$ is $\sigma$-cyclic. We refer to $\sigma$-cyclicity as shift-cyclicity and refer to $\sigma$-cyclic functions as shift-cyclic functions. Moreover, we write $S[f]$ instead of $\sigma[f]$, borrowing the notation from the $1$-variable case.
We end this section with a few properties of shift-cyclic functions that will be important for later discussion. Our first result is immediate from the definitions.
@@
(1.5) also implies that a shift-cyclic function $f$ is ‘almost invertible’. It is important to note, however, that having $1/f\in\mathcal{X}$ is not sufficient for the shift-cyclicity of $f$ (see [23, Theorem 1.4]). We shall study about this algebraic nature of shift-cyclicity in Section 4.2.2 below. Let us also mention another useful property that helps us determine the shift-cyclicity of polynomials.
@@
$\lim_{n\to\infty}\|P-P_{n}Pf\|_{\mathcal{X}}=\lim_{n\to\infty}\|P(1-P_{n}f)\|_{\mathcal{X}}=0.$
@@
The remaining assertion is immediate from the first part. ∎
@@ 2.1. On the unit disk $\mathbb{D}$
We begin by generalizing the space $H^{2}(\mathbb{D})$ introduced in Section 1.1. The motivation to study these spaces comes from a natural property of functions in $\operatorname{Hol}(\mathbb{D})$ that was discovered by Hardy [47].
For any $f\in\operatorname{Hol}(\mathbb{D})$, $r<1$ and $0<p<\infty$, we define the radial $p$-means
@@
By the maximum modulus principle, it is clear that $M_{\infty}(r,f)$ is non-decreasing in $r$ for a fixed $f$, and by the Hadamard three-circles theorem (see [94, Theorem 12.8]), it follows that $\log M_{\infty}(r,f)$ is a convex function of $\log r$. Hardy observed that the same is true for $M_{p}(r,f)$ for all values of $p$ [^2] Hardy mentioned in his paper that the case $p=1$ was suggested to him by H. Bohr (brother of the physicist N. Bohr) and E. Landau more than a year before he published his result..
Theorem 2.1 (Theorem I, [47]).
@@
In this section, we describe a few of the many different properties of Hardy space functions. We have cherry-picked only those that will be important to our discussion on shift-cyclicity. The reader is referred to [33, 39, 52, 83, 94] to learn more about the Hardy spaces. We start with the most fundamental of these properties. Let us write $\mathbb{T}:=\partial\mathbb{D}$ for the unit circle.
@@
References. This is essentially a result of Fatou [37], who proved a general result about radial limits of Poisson integrals of complex measures supported on $\mathbb{T}$. Fatou’s theorem gives the case $1\leq p\leq\infty$ while the rest follows from a result of F. and R. Nevanlinna [79], which states that for any $f\in H^{p}(\mathbb{D})$ with $0<p<\infty$, there exist $\varphi,\psi\in H^{\infty}(\mathbb{D})$ such that $f=\varphi/\psi$.
2.1.1. Boundary properties
For each $f\in H^{p}(\mathbb{D})$, we define its boundary function a.e. on $\mathbb{T}$ as
@@ 2.1.1. Boundary properties
Boundary functions play an important role in understanding the structure of functions in the Hardy spaces. Let us write $\mathcal{Z}(f)$ for the set of zeros in $\mathbb{D}$ of each $f\in\operatorname{Hol}(\mathbb{D})$ (counting multiplicities) and let $m(0)$ be the multiplicity of the zero at 00 (if $0\in\mathcal{Z}(f)$). Riesz [90] gave the following factorization result.
@@
$B(z)=z^{m(0)}\prod_{w\in\mathcal{Z}(f)\setminus\{0\}}\frac{|w|}{w}\frac{z-w}{1-\overline{w}z}.$
@@
Theorem 2.5 shows that
$\|f\|_{p}=\lim_{r\to 1^{-}}M_{p}(r,f),$
which is how $\|f\|_{p}$ is typically defined. It can then be checked that $\left\|\cdot\right\|_{p}$ is a norm on $H^{p}(\mathbb{D})$ for $1\leq p\leq\infty$, and that
@@
is a translation-invariant metric on $H^{p}(\mathbb{D})$ for $10<p<$ 1. Among other things, Riesz’s theorems allow us to obtain the following description of Hardy spaces. See [33, Section 3.2] for all the details.
@@ 2.1.2. Smirnov factorization and shift-cyclicity
    $f$ is said to be inner [^3] Beurling coined the terms inner/outer in [19]. Therefore, the terminology ‘Beurling inner/outer’ is quite common in the literature. if $|f|<1$ on $\mathbb{D}$ and $|f^{*}|=1$ a.e. on $\mathbb{T}$.
@@
    $f$ is said to be outer if $f(0)\neq 0$, $\log|f^{*}|\in L^{1}(\mathbb{T})$, and
    (2.2) $\log|f(0)|=\int_{0}^{2\pi}\log|f^{*}(e^{i\theta})|\frac{d\theta}{2\pi}.$
@@
Smirnov [101] and Beurling [19] used this factorization property to show that $f\in H^{2}(\mathbb{D})$ is shift-cyclic if and only if it is outer. In fact, Beurling showed that if we consider the inner part of $f$, i.e., $f_{inn}:=Bf_{\sigma}$, then
@@
Moreover, every closed shift-invariant subspace $E\subset H^{2}(\mathbb{D})$ is of the form $\varphi H^{2}(\mathbb{D})$ where $\varphi$ is the greatest common divisor of the inner factors of each $f\in E$ (see [52, Chapter 6 p. 85]), i.e.,
@@
Beurling’s approach extends to every $1\leq p<\infty$ using $L^{p}$-duality and the fact that $H^{p}(\mathbb{D})$ can be identified as a closed subspace of $L^{p}(\mathbb{T})$ (see [33, Theorem 7.4]). Lastly, note that we can also make sense of the shift-invariant subspace $S[f]$ for any $f\in H^{p}(\mathbb{D})$ with $10<p<$ 1 by declaring it as the closure of polynomial multiples of $f$. Gamelin [38] then showed that the Smirnov-Beurling theorem extends to the case $10<p<$ 1 as well. We summarize these results in the following theorem.
@@
This theorem was generalized by Lax [66] and Halmos [44] to the case of vector-valued Hardy spaces. Several other generalizations and modifications of this theorem exist that this article is too short to contain.